OpenStudy (stefrheart):

Write an equation for the line that is perpendicular to the line x-3y=9 at the x-intercept.

OpenStudy (aum):

First find the x-intercept. To do that, set y = 0 and solve for x. That will be the x-intercept.

OpenStudy (stefrheart):

okay so x=9

OpenStudy (aum):

Yes, the x-intercept is 9 and the point is (9,0) (because on the x-axis, which is where the x-intercept is, y = 0). Next, we need to find the slope of the line x-3y=9. Put it in the slope-intercept form y = mx + b so we can find the slope 'm'.

OpenStudy (aum):

x - 3y = 9 add 3y to both sides: x = 9 + 3y add -9 to both sides: x - 9 = 3y divide both sides by 3: 1/3 * x - 3 = y or y = 1/3 * x - 3 What is the slope of the line y = 1/3 * x - 3 ? Compare it to y = mx + b where 'm' is the slope.

OpenStudy (stefrheart):

the slope is 1/3. So do i replace the m with the 1/3 and the b with a 0?

OpenStudy (aum):

The slope of this line is 1/3. What will be the slope of a line that is perpendicular to the above line?

OpenStudy (stefrheart):

it would be -3

OpenStudy (aum):

Correct. So you know now the slope of the line is -3 and that it passes through the point (9,0) found earlier. What is the equation of this line?

OpenStudy (stefrheart):

it would be y=-3x?

OpenStudy (aum):

y = mx + b m = -3 y = -3x + b ---- (1) It passes through the point (9, 0). when x = 9, y = 0. Put it in (1). 0 = -3(9) + b 0 = -27 + b b = 27 y = -3x + 27 is the answer in slope-intercept form. 3x + y = 27 is the answer in standard form.

OpenStudy (stefrheart):

thank you so so so much!! :)

OpenStudy (aum):

You are welcome.